OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = 1/3x + 14/3 y = 3x + 10 y = -3x - 2 y = -1/3+ 10/3

OpenStudy (inkyvoyd):

<3

OpenStudy (texaschic101):

little hint : parallel lines will have the same slope

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown. line through the points 0 comma 6 and negative 2 comma 0 with a point 6 comma 2 not on the line y = 1/3x y = -1/3x + 4 y = 3x + 20

OpenStudy (anonymous):

So since we know parallel lines have the exact same slope, all but c is out. If you wanted to work it out, then substitute the x and y values from the "pass-through" point and solve for b.

OpenStudy (anonymous):

thankyew

OpenStudy (anonymous):

Not a problem :-)